# Gradient-free neural topology optimization

### v1

@@ 5 Discussion
To verify these limitations, we carry out a simple study on the expressivity of the model. We consider several sample designs (obtained with MMA) and pass them through the LBAE, to investigate whether the model has enough expressivity for reconstruction. As shown in Figure 7, the designs reconstructed from the latent encoding provided by the encoder are imperfect and lack the finer features. Next, we take the latent encoding of each design and optimize it (fitting MSE loss to the MMA design, using ADAM optimizer), to verify whether the model has enough expressive capacity. The results of the decoded designs with fitted latent vectors are shown in the bottom row. These designs incorporate finer features, providing more accurate and detailed representations of the MMA target designs. Nevertheless, the blurriness of these finer features is not completely eliminated, indicating expressivity limits of the decoder.
Figure 7: Examples of baseline designs obtained with MMA, and their reconstructions obtained with LBAE. Note that LBAE still exhibits some blurriness wwhen reconstructing finer features.
We speculate that these effects could be partially attributed to the convolutional layers – known to exhibit higher impedance to finer features [59, 60] and the spectral bias of neural networks [60]. In other words, convolutional neural nets struggle more with fitting fine features than coarse ones. This deficiency, however, could be potentially addressed by extending the training, or further adjustments in the architecture. As shown in the presented examples in Figure 3 lack of fine features can have a detrimental influence on the performance of the designs, and thus the overall optimization capabilities of the proposed method. Addressing those limitations might lead to significant improvements in the model’s performance.
Finally, the approach using variational autoencoder architectures based on convolutional neural networks is by design constrained to uniform meshes, and restricted in terms of resolution. Furthermore, we have not tried to solve 3D problems and we do not know how the approach would scale in those scenarios. Further research is required to explore new architectures that tackle these limitations.
@@
The second component of the performance of our method comes from the optimization procedure and its limitations. The major point of attention in that regard is the setting of the constraints. Specifically, the volume constraint was imposed following the method of [12], by applying a constrained sigmoid transformation on the output of the convolutional layers of the generator. However, this way of enforcing constraints could potentially hinder latent optimization. During the training, the VAE was exposed to samples with different volume fractions, which resulted not only in different shapes but in different topologies. Therefore, imposing the constraint directly in latent space by restricting to a manifold corresponding to a particular volume fraction could be more beneficial. Finding an alternative way of enforcing the constraint could potentially further improve the performance of the proposed latent optimization method.
@@
In our benchmarks, while comparing different methods, we used objective values reached at the same number of FEM simulations (simulation budget). This approach gives a fair overview of the total computational cost required for optimization, assuming that FEM comprises the majority of the computational cost in the optimization loop. However, most gradient-free optimizers (including CMA-ES used in this study) rely on evaluating a population (or a batch) of trial designs that can be done in parallel (lowering the total time significantly). This is not the case for gradient-based approaches, where the simulations are strictly sequential. In our set-up, the population sizes are in order 10, which means that the number of iterations (with parallelization – the compute time) is one order of magnitude less than the budget, as shown in the example in Figure 8. Consequently, while considering the parallelization, the gap to gradient-based approaches closes even further.
Figure 8: Comparison of loss curves for an example out of distribution problem. Given that for gradient-free optimization the simulations can be executed in parallel for each population, comparing performance based on iteration is more representative for the actual time (and not the computational cost, as with comparison by FEM evaluations). From this perspective, the gap between gradient-based and gradient-free optimizers closes even further.
@@ 6 Conclusion
In this study, we presented a proof of concept framework for optimizing topology designs without gradients in a latent space of a Latent-Bernoulli Autoencoder (LBAE). We found that the proposed method significantly improves the scalability of gradient-free optimization – addressing the major shortcoming of this approach. Our extensive benchmarks indicate that reparameterizing the optimization problem into a lower-dimensional latent space of LBAE leads to dramatic gains in performance, with at least one order of magnitude faster convergence as opposed to the conventional black-box optimization without reparameterization. We demonstrated the robustness of our method, by showing that the performance gains are significant even in the worst-case scenario i.e. considering only the worst runs out of all random initialization. Furthermore, we investigated the generalization performance, by deploying the LBAE model on a completely new problem (thermal compliance optimization) without retraining. Even though the design features desired in this problem are significantly different from those seen in the training dataset, the LBAE-driven optimization converges to a reasonable design, with performance close to that of the gradient-based baseline, proving promising generalization capabilities.
Nevertheless, our implementation still suffers from several limitations. We decided to impose hard constraints (volume fraction), which potentially hinders the optimization performance. We speculate, that imposing constraints directly in the latent space might potentially lead to better optimization. The results indicate a bias towards certain feature length scales in the designs. Generating features over a wider range of length scales, however, is desired for improving the expressivity and variety of the solutions. Finally, our current architecture is limited to uniform grid meshes and a restricted resolution. Convolutional layers may not scale well to 3D applications. Therefore, exploring different architectures such as Graph Neural Networks is expected to lead to even more significant performance gains, minimize the latent space dimension, further improve the scalability, and extend the applicability of latent space optimization. Still, we are surprised by the large performance improvements we found in our work relative to a conventional gradient-free approach to topology optimization, and we find particularly encouraging to show for the first time the generalization ability for out-of-distribution cases.
@@ Appendix A Complementary results
Figures 9, 10 and 11 present results for the remaining optimizer configurations: CMA-ES with the default settings, as well as CMA-ES and BIPOP-CMA-ES with customized settings (added variable clipping and initialization, both within the range of [-5, 5]).
Figure 9: Results for different reparameterization models, optimized with CMA-ES optimizer with default settings
Figure 10: Results for different reparameterization models, optimized with CMA-ES optimizer with customized settings
Figure 11: Results for different reparameterization models, optimized with BIPOP-CMA-ES optimizer with customized settings
@@ Appendix B Thresholded designs
Figures 12, 13 and 14 present the results for thresholded designs. The relative error is calculated with respect to the compliance obtained for a thresholded MMA design.
Figure 12: Statistics for the thresholded designs, optimized with CMA-ES with default settings.
Figure 13: Statistics for the thresholded designs, optimized with CMA-ES with custom settings (variable clipping and initialization within the range [-5, 5]).
Figure 14: Statistics for the thresholded designs, optimized with BIPOP-CMA-ES with custom settings (variable clipping and initialization within range [-5, 5]).
@@ B.1 Comparison of optimizers
Figures 15, 16 and 17 compare the performance of 2 different optimizers, with different settings (default and custom). In the custom configuration we apply clipping of the latent variables in the range of [-5, 5], and initialize the latent variable within that range.
Figure 15: Comparison of optimizers using VAE parameterization
Figure 16: Comparison of optimizers using conventional pixel parameterization
Figure 17: Comparison of optimizers using LBAE parameterization
@@ B.2 Optimization with thresholding
Figure 18: Caption
@@ B.3 Correction of the latent vector
To account for the limitations arising from imposing the volume constraint in a hard way, we experiment with a strategy to account for that effect by correcting the latent representation. Given a latent representation, we generate a target design with an imposed volume fraction constraint (sigmoid with constrained mean). Given the same latent representation, we generate a design without the constraint (applying only standard sigmoid activation in the output layer). The latent vector is then optimized by minimizing the MSE loss between the target design (with the volume constraint) and the design without the constraint. The gradients are backpropagated through the decoder network, and the updates to the latent vector are applied using ADAM optimizer, with learning rate 0.1, using 10 iterations. The results are compared with the standard approach, without correction in Figure 19.
@@ B.4 Convergence rate estimations
Figure 20 presents mean error (averaged over all designs from both datasets, disregarding 10% of the worst runs) vs number of objective evaluations. Although each problem is characterized by a different optimization history, affected e.g. by initialization, the plot gives a qualitative indication of the convergence rates between the 3 different parameterization schemes.
@@ Appendix C Shape similarity argument
Here we present a quantitative argument supporting our assumption that the design similarity could be a reasonable proxy for good performance. In other words, we show that designs optimized for different objectives oftentimes are similar (share the same features). We consider the designs of the MBB-beam reported in work by Desai et al. [4], and compare the design optimized for compliance against designs optimized for fracture with different hyperparameter settings. We quantify the similarity using the Learned Perceptual Image Patch Similarity (LPIPS) metric [45], which provides superior performance for quantifying perceptual similarity by using the feature maps of pre-trained neural networks. The similarity indices for different designs are shown in Figure 21, computed using two different feature map extractors (VGG and Alex-Net).
The compliance design is slightly more dissimilar to the fracture designs - as indicated by a higher LPIPS index. However, we also find a counter-example - the fracture design (extracted from figure 8b in [4]), with a comparably high similarity index. In other words, the fracture design (8b) is as similar to the remaining fracture designs as the compliance design. The findings are consistent regardless of which feature extractor was used for LPIPS calculation.
@@ Appendix D Latent dimensionality and population size trade-off
Table 1 illustrates the trade-off present while choosing the latent space dimensionality. Models with larger latent dimensionality are easier to train (requiring less compression). On the other hand, larger dimensionality hinders the optimization performance of gradient-free optimizers. In our study, the latent dimensionality of 256 was found to balance these two factors best.
@@
Similarly, the optimal population size (number of samples for an update of the gradient-free optimizer), is also governed by a trade-off. In this case, it is the frequency of the updates (lower population – more frequent updates) against the quality of the updates (higher population – better estimation of the optimal update). In that regard, the population size on the order of 10 samples provides the best performance.
Appendix E Alternative configurations of the baseline VAE model

### v2

@@ 5 Discussion
To verify these limitations, we performed a simple study on the expressivity of the model. We consider several sample designs (obtained with MMA) and pass them through the LBAE, to investigate whether the model has enough expressivity for reconstruction. As shown in Figure 11, the designs reconstructed from the latent encoding provided by the encoder are imperfect and lack the finer features. Next, we take the latent encoding of each design and optimize it (fitting MSE loss to the MMA design, using ADAM optimizer), to verify whether the model has enough expressive capacity. The results of the decoded designs with fitted latent vectors are shown in the bottom row. These designs incorporate finer features, providing more accurate and detailed representations of the MMA target designs. Nevertheless, the blurriness of these finer features is not eliminated, indicating the expressivity limits of the decoder.
Figure 11: Examples of baseline designs obtained with MMA, and their reconstructions obtained with LBAE. Note that LBAE still exhibits some blurriness when reconstructing finer features.
We speculate that these effects could be partially attributed to the convolutional layers – known to exhibit higher impedance to finer features [82, 83] and the spectral bias of neural networks [83]. In other words, convolutional neural nets struggle more with fitting fine features than coarse ones. This deficiency, however, could be potentially addressed by extending the training, or further adjustments in the architecture. As shown in the examples presented in Figure 4, a lack of fine features can have a detrimental influence on the performance of the designs and thus the overall optimization capabilities of the proposed method. Addressing those limitations might lead to significant improvements in the model’s performance.
Finally, the approach using variational autoencoder architectures based on convolutional neural networks is by design constrained to uniform meshes, and restricted in terms of resolution. Furthermore, we have not tried to solve 3D problems, and we do not know how the approach would scale in those scenarios.
@@
The second component of the performance of our method comes from the optimization procedure and its limitations. The major point of attention in this regard is the setting of the constraints. Specifically, the volume constraint was imposed following the method of [19], by applying a constrained sigmoid transformation on the output of the convolutional layers of the generator. However, this way of enforcing constraints could potentially hinder latent optimization. During training, the VAE was exposed to samples with different volume fractions, resulting in not only different shapes but also different topologies. Therefore, imposing the constraint directly in latent space by restricting to a manifold corresponding to a particular volume fraction could be more beneficial. Alternatively, the volume fraction could be an input to the model. In this way, the model would be trained to output designs that satisfy (approximately) the volume constraint. Finding a different way of enforcing the constraint could potentially further improve the performance of the proposed latent optimization method.
@@
In our benchmarks, we compare different methods based on the budget of objective function evaluations (number of simulations) and optimizer iterations. In a single iteration, gradient-free optimizers usually evaluate a whole population of trial solutions, i.e. a simulation is run for each trial solution. However, these simulations are run in parallel, which means that the iteration time remains comparable to the compute time of a single simulation (assuming that the FEM simulation comprises the majority of the computational cost in the optimization loop). In other words, the budget (number of simulations) is more representative of the computational resources (CPU time), while the number of iterations is more representative of the absolute compute time. An example comparison between different methods using both metrics is shown in Figure 12. Since the population size in our set-up is on the order of 10, the difference between budget and iterations for gradient-free approaches is approximately 1 order of magnitude. In the gradient-based approach, there is a single simulation per iteration (although in practice, in general, the cost of computing sensitivities can be comparable to the cost of the simulation).
Figure 12: Comparison of loss curves for an example out of distribution problem. Given that for gradient-free optimization the simulations can be executed in parallel for each population, comparing performance based on iteration is more representative of the actual time (and not the computational cost, as with comparison by FEM evaluations). From this perspective, the gap between gradient-based and gradient-free optimizers closes even further.
Although our approach significantly closes the gap between gradient-free optimization without parameterization and gradient-based methods, as shown in Figure 12, it still remains relatively expensive in terms of computational resources for smooth and differentiable problems. We do not expect this approach to be competitive on problems that can be solved efficiently with gradient-based methods. However, with our validation example, we show that our method enables gradient-free optimization with reasonable computational resources and even achieves better objective values for problems where gradient-based algorithms struggle.
Finally, our validation example is relatively limited. It can be argued that the issues of the gradient-based optimizer can be overcome with a number of techniques such as e.g. continuation schemes, hyperparameter tuning, or alternative formulations of the optimization problems [15, 9]. Alternatively, these challenges could be tackled with an emerging approach that involves the use of Physics-Informed Gaussian Processes, which optimizes the solution of the governing PDEs simultaneously with the design variables, thereby eliminating the need for continuation schemes [59]. While with these techniques better performance may be reached, we also want to highlight the robustness of our method to the choice of hyperparameters and initialization – an important feature from the practitioner’s perspective. Nevertheless, future work is encouraged to further assess the merits of our method.
@@ 6 Conclusion
In this study, we presented a framework for optimizing topology designs without gradients in a latent space of a Latent-Bernoulli Autoencoder (LBAE). We found that the proposed method significantly improves the scalability of gradient-free optimization – addressing the major shortcoming of this approach.
Our extensive benchmarks indicate that reparameterizing the optimization problem into a lower-dimensional latent space of LBAE leads to dramatic gains in performance, with at least one order of magnitude faster convergence as opposed to the conventional black-box optimization without reparameterization. We demonstrated the robustness of our method, by showing that the performance gains are significant even in the worst-case scenario i.e. considering only the worst runs out of all random initialization.
We applied our method without retraining to optimize a structure undergoing brittle fracture and demonstrated that it is capable of outperforming the gradient-based approach on problems characterized by discontinuous loss landscapes, where gradient-based optimizers tend to struggle. Specifically, our proposed method outperformed the traditional gradient-based optimizer (MMA) by a significant margin, consistently delivering an improvement of the order of 30% on all tested configurations with different initialization strategies and hyperparameter settings.
Nevertheless, our implementation still suffers from several limitations. We decided to impose hard constraints (volume fraction), which potentially hinder the optimization performance. We speculate that imposing constraints directly in the latent space might potentially lead to better optimization. The results indicate a bias towards certain feature length scales in the designs. Generating features over a wider range of length scales is desired to improve the expressivity and variety of the solutions. Our current architecture is limited to uniform grid meshes and a restricted resolution. Convolutional layers may not scale well to 3D applications. Therefore, exploring different architectures is expected to lead to even more significant performance gains, minimize the dimension of the latent space, further improve the scalability, and extend the applicability of latent space optimization. Finally, we limited our study to density-based optimization and volume fraction constraint. Exploring other parameterization approaches, such as level-set method, and other types of constraints, remains a subject for future work.
Still, we are surprised by the large performance improvements we found in our work relative to a conventional gradient-free approach to topology optimization, and we find it particularly encouraging to show for the first time the generalization ability on problems, where the traditional gradient-based optimizers reach their limitations.
@@ Appendix A Complementary results
Figures 13, 14, and 15 present results for the remaining optimizer configurations: CMA-ES with default settings, as well as CMA-ES and BIPOP-CMA-ES with customized settings (added variable clipping and initialization, both within the range of [-5, 5]).
Figure 13: Results for different reparameterization models, optimized with CMA-ES optimizer with default settings.
Figure 14: Results for different reparameterization models, optimized with CMA-ES optimizer with customized settings.
Figure 15: Results for different reparameterization models, optimized with BIPOP-CMA-ES optimizer with customized settings.
@@ Appendix B Thresholded designs
Figures 16, 17, and 18 present the results for thresholded designs. The relative error is calculated with respect to the compliance obtained for a thresholded MMA design.
Figure 16: Performance of the thresholded designs, optimized with CMA-ES with default settings.
Figure 17: Performance of the thresholded designs, optimized with CMA-ES with custom settings (variable clipping and initialization within the range [-5, 5]).
Figure 18: Performance of the thresholded designs, optimized with BIPOP-CMA-ES with custom settings (variable clipping and initialization within the range [-5, 5]).
@@ B.1 Comparison of optimizers
Figures 19, 20, and 21 compare the performance of two different optimizers, with different settings (default and custom). In the custom configuration, we apply clipping of the latent variables in the range of [-5, 5], and initialize the latent variable within that range.
Figure 19: Comparison of optimizers using VAE parameterization.
Figure 20: Comparison of optimizers using conventional pixel parameterization.
Figure 21: Comparison of optimizers using LBAE parameterization.
@@ B.2 Optimization with thresholding
Figure 22: Performance of designs optimized with CMA-ES, using thresholding during optimization.
@@ B.3 Correction of the latent vector
To account for the limitations arising from imposing the volume constraint in a hard way, we experiment with a strategy to account for that effect by correcting the latent representation. Given a latent representation, we generate a target design with an imposed volume fraction constraint (sigmoid with a constrained mean). Given the same latent representation, we generate a design without the constraint (applying only standard sigmoid activation in the output layer). The latent vector is then optimized by minimizing the MSE loss between the target design (with the volume constraint) and the design without the constraint. Gradients are backpropagated through the decoder network, and updates to the latent vector are applied using the ADAM optimizer, with a learning rate of 0.1, using 10 iterations. The results are compared with the standard approach, without correction in Figure 23.
@@ B.4 Convergence rate estimations
Figure 24 presents the mean error (averaged over all designs from both datasets, disregarding 10% of the worst runs) vs the number of objective evaluations. Although each problem is characterized by a different optimization history, affected e.g. by initialization, the plot gives a qualitative indication of the convergence rates between the 3 different parameterization schemes.
@@ Appendix C Shape similarity argument
Here, we present a quantitative argument supporting our assumption that design similarity could be a reasonable proxy for good performance. In other words, we show that designs optimized for different objectives oftentimes are similar (share the same features). We consider the designs of the cantilever reported in the work by Desai et al. [9], and compare the design optimized for compliance with the designs optimized for fracture with different hyperparameter settings. We quantify similarity using the Learned Perceptual Image Patch Similarity (LPIPS) metric [61], which provides superior performance to quantify perceptual similarity using the feature maps of pre-trained neural networks. The similarity indices for different designs are shown in Figure 25, calculated using two different feature map extractors (VGG and Alex-Net).
The compliance design is slightly more dissimilar to the fracture designs - as indicated by a higher LPIPS index. However, we also find a counterexample, the fracture design (extracted from Figure 8b in [9]), with a comparably high similarity index. In other words, the fracture design (8b) is as similar to the remaining fracture designs as the compliance design. The findings are consistent regardless of which feature extractor was used for the LPIPS calculation.
@@ Appendix D Latent dimensionality and population size trade-off
Table 1 illustrates the trade-off present while choosing the dimensionality of the latent space. Models with larger latent dimensionality are easier to train (requiring less compression). On the other hand, larger dimensionality hinders the optimization performance of gradient-free optimizers. In our study, the latent dimensionality of 256 was found to balance these two factors best.
@@
Similarly, the optimal population size (number of samples for an update of the gradient-free optimizer), is also governed by a trade-off. In this case, it is the frequency of the updates (lower population – more frequent updates) compared to the quality of the updates (higher population – better estimation of the optimal update). In that regard, the population size on the order of 10 samples provides the best performance.
Appendix E Model Architectures
The details of the LBAE model architectures are listed in Table 2 and 3.
Table 2: LBAE Encoder Architecture. In each residual block, the residual connection is defined between the output of the first and last batch norm.
Input features block
Conv2d(1, 64, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
LeakyReLU(negative_slope=0.2)
Residual Blocks (3)
Block 1
Conv2d(64, 64, kernel_size=(4, 4), stride=(2, 2), padding=(1, 1), bias=False)
BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
Conv2d(64, 64, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
Conv2d(64, 64, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
LeakyReLU(negative_slope=0.2)
Block 2
Conv2d(64, 128, kernel_size=(4, 4), stride=(2, 2), padding=(1, 1), bias=False)
BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
Conv2d(128, 128, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
Conv2d(128, 128, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
LeakyReLU(negative_slope=0.2)
Block 3
Conv2d(128, 256, kernel_size=(4, 4), stride=(2, 2), padding=(1, 1), bias=False)
BatchNorm2d(256, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
Conv2d(256, 256, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(256, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
Conv2d(256, 256, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(256, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
Latent features block
Conv2d(256, 256, kernel_size=(4, 4), stride=(2, 2), padding=(1, 1), bias=False)
LeakyReLU(negative_slope=0.2)
Linear(in_features=4096, out_features=256, bias=True)
Tanh()
Table 3: LBAE Decoder Architecture. In each residual block, the residual connection is defined between the output of the first and last batch norm.
Latent features block
Linear(in_features=256, out_features=4096, bias=True)
LeakyReLU(negative_slope=0.2)
Residual Blocks (3)
Block 1
ConvTranspose2d(256, 256, kernel_size=(4, 4), stride=(2, 2), padding=(1, 1), bias=False)
BatchNorm2d(256, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
ConvTranspose2d(256, 256, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(256, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
ConvTranspose2d(256, 256, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(256, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
LeakyReLU(negative_slope=0.2)
Block 2
ConvTranspose2d(256, 128, kernel_size=(4, 4), stride=(2, 2), padding=(1, 1), bias=False)
BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
ConvTranspose2d(128, 128, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
ConvTranspose2d(128, 128, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
LeakyReLU(negative_slope=0.2)
Block 3
ConvTranspose2d(128, 64, kernel_size=(4, 4), stride=(2, 2), padding=(1, 1), bias=False)
BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
ConvTranspose2d(64, 64, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
ConvTranspose2d(64, 64, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
LeakyReLU(negative_slope=0.2)
Output block
ConvTranspose2d(64, 64, kernel_size=(4, 4), stride=(2, 2), padding=(1, 1), bias=False)
BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
ConvTranspose2d(64, 1, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), bias=False)
E.1 Alternative configurations of the baseline VAE model